OpenStudy (anonymous):

MEDAL GIVEN TO CORRECT ANSWER. A book with mass 2.3 kg sits on a table. What is the normal force on the book? A. 22.5 N B. 2.3 N C. 18.3 N D. 9.8 N

OpenStudy (anonymous):

\[\sum(Fy)\]= 0 (because the system is in equilibrium) = N- mg N=mg = 2.3*(9.8)= 22.5 N The correct answer is A.